OpenStudy (anonymous):

The grades on the last math exam had a mean of 88%. Assume the population of grades on math exams is known to be normally distributed with a standard deviation of 6. What percent of students earn a score between 72% and 90%?

ganeshie8 (ganeshie8):

start by finding the zscores for both observations

OpenStudy (anonymous):

How? My lessons are not very explanitory.

ganeshie8 (ganeshie8):

\(\large Zscore = \frac{\text{observation - mean}}{\text{standard deviation}}\)

ganeshie8 (ganeshie8):

you're given : mean = 88 standard deviation = 6

OpenStudy (anonymous):

Ok, so 14.6?

ganeshie8 (ganeshie8):

Zscore for 72 = \(\large \frac{72-88}{6} = ?\) Zscore for 90 = \(\large \frac{90-88}{6} = ?\)

OpenStudy (anonymous):

-2, and 0.33?

ganeshie8 (ganeshie8):

you should get : Zscore for 72 = -2.67 Zscore for 90 = 0.33

ganeshie8 (ganeshie8):

next, find the corresponding areas

ganeshie8 (ganeshie8):

knw how to grab areas for zscore from zscore table ?

OpenStudy (anonymous):

Assume that I do not know how to do any of this, because I don't.

ganeshie8 (ganeshie8):

okay, its going to be very easy

ganeshie8 (ganeshie8):

http://www.math.upenn.edu/~chhays/zscoretable.pdf

ganeshie8 (ganeshie8):

^^thats the zscore table

ganeshie8 (ganeshie8):

area for zscore of -2.67 = \(0.0038\) area of zscore of 0.33 = \(0.6293\)

ganeshie8 (ganeshie8):

see if u can figure out how i got above areas from zscore table

ganeshie8 (ganeshie8):

So, the percent of students earn a score between 72% and 90% = 0.6293 - 0.0038 = 0.6255 = 62.55%

ganeshie8 (ganeshie8):

let me knw if smthng doesnt make sense...

OpenStudy (anonymous):

I tried so hard to understand that table in my lesson, and honestly It looks like spanish.....

ganeshie8 (ganeshie8):

hahah let me explain u quick

ganeshie8 (ganeshie8):

like any table, it has rows and columns

OpenStudy (anonymous):

Yes?

ganeshie8 (ganeshie8):

\( \begin{array}{|c|c|} \hline z&0.09&0.08&0.07&0.06&0.05&0.04&0.03&0.02&0.01&0.00& \\ \hline -3.4&0.0002&0.0003&0.0003&0.0003&0.0003&0.0003&0.0003&0.0003&0.0003&0.0003& \\ \hline -3.3&0.0003&0.0004&0.0004&0.0004&0.0004&0.0004&0.0004&0.0005&0.0005&0.0005& \\ \hline \end{array} \)

ganeshie8 (ganeshie8):

interpreting it is very easy : suppose your zscore is \(-3.45\)

ganeshie8 (ganeshie8):

then, put ur thumb on the row starting with \(-3.4\) and put another thumb on the column with \(0.05\)

ganeshie8 (ganeshie8):

\( \begin{array}{|c|c|} \hline z&0.09&0.08&0.07&0.06&\color{red}{0.05}&0.04&0.03&0.02&0.01&0.00& \\ \hline \color{Red}{-3.4}&0.0002&0.0003&0.0003&0.0003&0.0003&0.0003&0.0003&0.0003&0.0003&0.0003& \\ \hline -3.3&0.0003&0.0004&0.0004&0.0004&0.0004&0.0004&0.0004&0.0005&0.0005&0.0005& \\ \hline \end{array} \)

ganeshie8 (ganeshie8):

you wud get the area : \(0.0003\)

ganeshie8 (ganeshie8):

so for zscore of \(-3.45\), the area is \(0.0003\)

ganeshie8 (ganeshie8):

see if that makes more or less sense...

OpenStudy (anonymous):

More sense haha thank you

ganeshie8 (ganeshie8):

good :) may i ask u to find the area for zscore of \(-3.38\) ?

OpenStudy (anonymous):

0.004?

ganeshie8 (ganeshie8):

you got it !!

ganeshie8 (ganeshie8):

and did u get why we subtracted both areas and concluded answer is 62.55% ?